Displaying report 1-1 of 1.
Reports until 17:23, Tuesday 03 September 2024
Controls General
Jeffrey Wack - posted 17:23, Tuesday 03 September 2024 (11856)
775 Piezo Lock on Cavity #2

[Jeff, Torrey]

Connected the large ThorLabs piezo mounted on cavity #2 directly to the Moku (no HV amplifier).

Transmission voltage -10.7mV (no light) -9.5mV (on resonance). The photodetector in transmission was selected to measure ringdowns, it is not rated for light below 800nm. Responsivity in amps/Watt is less than 0.1

Max error amplitute is 4mV pkpk, and locking with the piezo results in constant flashing, saturating the error signal.

What voltages are we giving the piezo? The Moku has output limits of 10V. Since the maximum amplitude of the error signal is 2mV, the maximum gain we can have in our controller before saturating the Moku output is 20log_10(10/0.002) = 74dB. This is definitely being exceeded at low frequencies, we will try the HV amplifier.

The piezo for cavity control is now routed through the HV amplifier. We are able to achieve better locks but still no where sufficient for stable cavity operation. The error signal is being saturated in both directions, same as before. After some trouble shooting, two things are definitively contributing to this poor lock. The fans being on causes a non-trivial amount of noise for the piezo to handle. Additionally we found that the 840 Hz known to be coming from the poor grounding on the DC modulation port of the laser is interfering somehow. With the HV source in play we can scan and lock fully on the piezo. Getting rid of the fans and this noise introduced by having a cable plugged in into the laser seems to give a quality of lock similar to that of locking only with the laser.

 

 

Images attached to this report
Displaying report 1-1 of 1.